OpenStudy (anonymous):

What is the equation of a line passing through (-8, -3) and having a slope of -4?

OpenStudy (anonymous):

y = −4x + 20 y = 4x + 20 y = −4x – 35 y = 4x – 35 y = -4x + 4

OpenStudy (marihelenh):

Use the equation, \[(y-y _{1})=m(x-x _{1})\] m=slope.

OpenStudy (anonymous):

ok so it's not the first 2 right

OpenStudy (marihelenh):

Nope

OpenStudy (anonymous):

it's the 3rd one

OpenStudy (marihelenh):

Nor the last one either. \[(y-(-3))=-4(x-(-8))\]

OpenStudy (marihelenh):

Yes it is! I guess I didn't need to write that out. Good job!

OpenStudy (anonymous):

thank u